queen999:

Kelly tells you that when variables are in the denominator, the equation one over two plus three over x equals three over four becomes unsolvable. Kelly explains, "There is a value for x that makes the denominator zero, and you can't divide by zero." Demonstrate to Kelly how the equation is still solvable and explain your reasoning

supie:

\[1/2 + 3/x = 3/4\]\[3/x = 3/4 - 1/2\]\[3/x = 3/4 - 2/4\] (Now we Cross Multiply)\[(1)(x) = (3)(4)\] So\[ x = 12\]